# CMB observables and reheat temperature as a window to models of inflation and freeze-in dark matter production

### v1

@@ 1 Introduction
The standard paradigm is that, in its early evolution, the universe underwent several distinct phases. Its accelerated expansion (inflation) was followed by a period when the energy stored in the inflaton field was converted to relativistic particles, broadly referred to as a reheating period, during the oscillations of the inflaton field around the minimum of its potential. After that, the standard radiation-dominated (and later, matter-dominated) evolution of the universe took place. It has been emphasized for a long time that the CMB data give us important insight into the inflationary and reheating periods, constraining theoretical models of both. In this paper we propose a systematic approach for using CMB data for discriminating between different models of inflation and using the reheating temperature as a portal to constraining particle physics scenarios that are sensitive to it. As an example we discuss certain freeze-in DM scenarios.
@@
There are striking consequences of this fact. The model independent bounds on the reheating temperature, the BBN lower bound, arising due to the precision measurements of primordial abundance of light nuclei, and the upper bound of the order of the GUT/Planck scale [11], translate themselves, for each class of models, into very narrow ranges of the spectral index $n_{s}(k)$ and with interesting dependence on the parameter $r$. That is a strong test of inflationary models by the present and future CMB data. Moreover, one can investigate the implications for models of inflation of imposing some constraints on the reheating temperature motivated by various physical arguments. For instance, successful standard thermal leptogenesis  [12] requires $T_{\rm x}\geq 10^{9}$ GeV [13]. [^2] It was shown in ref. [14] that in non-thermal leptogenesis during the reheating period, CMB observables could probe the scales of seesaw and leptogenesis. A different, contradictory, example are upper bounds on reheat temperature derived when solving the so-called gravitino problem in certain models of supersymmetry breaking in supergravity with $T_{\rm x}\leq 10^{7}$ GeV [15, 16, 17, 18, 19], and axion and axion-like production as DM during reheating [20, 21, 22, 23].
@@ 2 CMB, inflaton potentials and reheating temperature
where $\Lambda_{\rm inf}$ represents a mass scale that determines the energy scale of the inflation and $\sqrt{\alpha}M_{\rm P}$ is an effective scale that can be higher than $M_{\rm P}$. A special case of the E-model, with $\alpha=1$ and $n=1$, mimics the standard Higgs-Starobinsky inflaton potential [4].
@@
The exponent $2n$ in the potential (2.4) is an even integer and we consider it as a number which defines the model. The other parameters $\Lambda_{\rm inf},\alpha$ and the inflaton field value $\phi_{k}$ when the mode with the co-moving wavenumber $k$ left the horizon can be expressed in terms of the three observables by the inverse relations. It is convenient to define the following combination of the CMB observables, $n_{s}$ and $r$, and the integer potential parameter $n$:
@@
Since at present only the upper bound on $r$ is known experimentally, it is interesting to illustrate the above results in the plane ($n_{s}$, $r$) as a function $r(n_{s})=f(n,A_{s},n_{s},T_{\times})$, for several fixed values of the exponent $n$. First, we compare our approach to the selection of the acceptable inflationary models based on the reference to the reheating temperature $T_{\times}$ with the one using an ansatz for the e-fold number $N_{k}$ to be in the range $50\div 60$. In figure 2 two sets of curves in the ($n_{s}$, $r$) plane are compared: colored curves with fixed $T_{\times}$, $r(n_{s})=f(n,A_{s},n_{s},T_{\times})$, and black ones with fixed $N_{k}$, $r(n_{s})=f(n,A_{s},n_{s},N_{k})$. The characters of those two sets of curves are quite different, especially for smaller values of $r$. A given model may be self-consistent only if it predicts a point on the ($n_{s}$, $r$) plane between the outermost colored curves because only then the consistency condition (2.17) and the model independent bound on the reheat temperature (2.35) are fulfilled. [^9] The highest $T_{\times}$ presented in figures in this section equals $1.6\cdot 10^{15}$ GeV and is somewhat smaller than the model-independent upper bound (2.35). The corresponding curves in figure 2 are quite short because the consistency condition (2.17) with so high $T_{\times}$ have solutions only for relatively big values of $r$. For $T_{\times}$ approaching the upper bound the length of the corresponding curve shrinks to a point. In many models the allowed region in the ($n_{s}$, $r$) plane has only some small overlap with the region corresponding to the often used condition $50\leq N_{k}\leq 60$. Clearly, the approach based on well motivated bounds on $T_{\times}$ may give significantly different results compared to an (essentially arbitrary) ansatz for $N_{k}$. Thus, in the following the dependence of the results on the reheating temperature (and not on $N_{k}$) will be investigated.
@@
Results for $r\geq 10^{-4}$ are presented in figures 3 and 4 where curves of constant $T_{\times}$ are shown together with present experimental bounds and sensitivity reaches of future experiments. The present $2\sigma$ contours of the Planck+BK18 data differ quite strongly from the global fit including also ACT and DESI data. The latter fit is more difficult to reconcile with the considered models and shows some preference for P-models and for E and T-models with large values of $n$. Actually, we see that in the considered models there is the upper bound on $n_{s}$ around 0.970, to be compared with the $2\sigma$ upper bound of the global fit at 0.980. One can also ask how constraining are the present CMB data for the reheating temperature. Given the large difference between the Planck data and the global fit one concludes that the whole range of temperatures is acceptable. Taken Planck and the global fit separately, some constraints on the reheating temperatures can be seen in the figures. Furthermore, it is worth noting the change of ordering of the $T_{\times}={\rm const}$ curves from small to large values, as a function of $n_{s}$. This is related to the fact that $w<\frac{1}{3}$ for $n=1$ while $w>\frac{1}{3}$ for $n>2$. The case of $n=2$ is a special one, where $w=\frac{1}{3}$ and the function $r(n_{s})$ does not depend on $T_{\times}$ (the black curve in figure 3). Improving the precision of $n_{s}$ measurements, as expected for the CMB S4 experiment, would be a stronger test of the models. The expected $2\sigma$ range (0.964, 0.974) would be particularly constraining if, in addition, the upper bound on $r$ was pushed down, say to $r<10^{-5}$ (see figure 2). It also follows from figure 2 that, if the present $2\sigma$ lower bound is valid for any value of $r$, then E- and T-models with $n=1$ predict a lower bound on $r$ of the order of $10^{-10}$. The corresponding bounds for P-models and all considered models with big values of $n$ are much weaker (e.g. of the order of $10^{-26}$ for $n=10$). Some differences in the predictions for $r(n_{s})$ for a fixed reheating temperature $T_{\times}$, particularly between P-model compared to E- and T-models, can be traced back to the shape of the inflaton potential in those models.
@@ 3 $T_{\times}$ sensitive dark matter production
Many freeze-in models with the DM production described by dim 5 and dim 6 effective operators have been discussed in the literature. They include, for instance, two scalar thermal bath particles $\Phi$ annihilating into a pair of DM fermions $\chi$, $\mathcal{L}=|\Phi|^{2}\chi\chi/\Lambda_{\rm DM}$, or two thermal bath fermions $\Psi$ annihilating either into a pair of scalar DM particles $S$, ${\cal L}={\bar{\Psi}}\Psi SS/\Lambda_{\rm DM}$ or into a pair of DM fermions $\chi$, ${\cal L}={\bar{\Psi}}\Psi{\bar{\chi}}\chi/\Lambda_{\rm DM}^{2}$.
@@ 4 Conclusions
The present and future experimental programmes will explore the properties of the Cosmic Microwave Background in a great detail. In this paper we have proposed a systematic approach to use the CMB observables and the model-independent bounds on the reheating temperature a) for discriminating between different models of inflation, b) to constrain a reheating temperature dependent particle physics scenario, freeze-in DM production models. In the first step of this approach, all independent parameters of an inflationary model are expressed in terms of the CMB observables. Next, adopting a one parameter effective description (by an effective dissipation rate $\Gamma$), a consistency relation fixes the reheating temperature in terms of the measured values of the CMB observables. For a given model of inflation the bounds on those observables then follow from model-independent bounds on the reheating temperature (see section 2). The striking consequence of those bounds is that a given model of inflation gives very narrow ranges of the power spectrum index $n_{s}$, different for different inflationary models and with very interesting dependence on the value of $r$. In this paper, the above described strategy has been applied to $\alpha$-attractor inflaton potentials, with three free parameters. For all the models we obtain an upper bound on the value of the spectral index, $n_{s}\lesssim 0.970$, making them more easily compatible with the Planck data than with the ACT data (see figures 3 and 4). We stress that those results are obtained for a very simple reheating model, which strongly correlates the effective equation of state parameter $w$ during the reheating period with the inflaton potential (see eq. (2.22)) .
@@ Acknowledgments
Authors thank S K Jeesun for comments. P.K. thanks Adeela Afzal and Shiladitya Porey for sharing their Mathematica codes used to produce the experimental contours in section 2.
@@ Appendix B Simplifying assumptions underlying this analysis
  • In the calculations regarding DM we have assumed instant and local thermal equilibrium to have been reached during the reheating period.

### v2

@@ 1 Introduction
The standard paradigm is that, in its early evolution, the universe underwent several distinct phases. Its accelerated expansion (inflation) was followed by a period when the energy stored in the inflaton field was converted to relativistic particles, broadly referred to as a reheating period, during the oscillations of the inflaton field around the minimum of its potential. After that, the standard radiation-dominated (and later, matter-dominated) evolution of the universe took place. It has been emphasized for a long time that the CMB data give us important insight into the inflationary and reheating periods, constraining theoretical models of both [1]. In this paper we propose a systematic approach for using CMB data for discriminating between different models of inflation and using the reheating temperature as a portal to constraining particle physics scenarios that are sensitive to it. As an example we discuss certain freeze-in DM scenarios.
@@
There are striking consequences of this fact. The model independent bounds on the reheating temperature, the BBN lower bound, arising due to the precision measurements of primordial abundance of light nuclei, and the upper bound of the order of the GUT/Planck scale [15], translate themselves, for each class of models, into very narrow ranges of the spectral index $n_{s}(k)$ and with interesting dependence on the parameter $r$. That is a strong test of inflationary models by the present and future CMB data. Moreover, one can investigate the implications for models of inflation of imposing some constraints on the reheating temperature motivated by various physical arguments. For instance, successful standard thermal leptogenesis  [16] requires $T_{\rm x}\geq 10^{9}$ GeV [17]. [^3] It was shown in ref. [18] that in non-thermal leptogenesis during the reheating period, CMB observables could probe the scales of seesaw and leptogenesis. A different, contradictory, example are upper bounds on reheat temperature derived when solving the so-called gravitino problem in certain models of supersymmetry breaking in supergravity with $T_{\rm x}\leq 10^{7}$ GeV [19, 20, 21, 22, 23, 24], and axion and axion-like production as DM during reheating [25, 26, 27, 28].
@@ 2 CMB, inflaton potentials and reheating temperature
where $\Lambda_{\rm inf}$ represents a mass scale that determines the energy scale of the inflation and $\sqrt{\alpha}M_{\rm P}$ is an effective scale that can be higher than $M_{\rm P}$. [^5] We use the same normalization of $\alpha$ in all three models. Observe that it is different from the normalization for T-model used e.g. in [33]. A special case of the E-model, with $\alpha=1$ and $n=1$, mimics the standard Higgs-Starobinsky inflaton potential [5].
@@
The exponent $2n$ in the potential (2.4) is an even integer and we consider it as a number which defines the model. [^6] We thank Renata Kallosh and Andrei Linde for drawing our attention to the fact that for polynomial potentials like P-model fractional values of $n$ are also of theoretical interest. The other parameters $\Lambda_{\rm inf},\alpha$ and the inflaton field value $\phi_{k}$ when the mode with the co-moving wavenumber $k$ left the horizon can be expressed in terms of the three observables by the inverse relations. It is convenient to define the following combination of the CMB observables, $n_{s}$ and $r$, and the integer potential parameter $n$:
@@
Since at present only the upper bound on $r$ is known experimentally, it is interesting to illustrate the above results in the plane ($n_{s}$, $r$) as a function $r(n_{s})=f(n,A_{s},n_{s},T_{\times})$, for several fixed values of the exponent $n$. First, we compare our approach to the selection of the acceptable inflationary models based on the reference to the reheating temperature $T_{\times}$ with the one using an ansatz for the e-fold number $N_{k}$ to be in the range $50\div 60$. In figure 2 two sets of curves in the ($n_{s}$, $r$) plane are compared: colored curves with fixed $T_{\times}$, $r(n_{s})=f(n,A_{s},n_{s},T_{\times})$, and black ones with fixed $N_{k}$, $r(n_{s})=f(n,A_{s},n_{s},N_{k})$. For $r\gtrsim 10^{-3}$ the two sets of curves have similar $r$-dependence, with the fixed temperature curves consistent with some fixed $N_{k}$, but not always in the range $50\div 60$. The characters of those two sets of curves as functions of $r$ are quite different for smaller values of $r$. A given model may be self-consistent only if it predicts a point on the ($n_{s}$, $r$) plane between the outermost colored curves because only then the consistency condition (2.17) and the model independent bound on the reheat temperature (2.35) are fulfilled. [^12] The highest $T_{\times}$ presented in figures in this section equals $1.6\cdot 10^{15}$ GeV and is somewhat smaller than the model-independent upper bound (2.35). The corresponding curves in figure 2 are quite short because the consistency condition (2.17) with so high $T_{\times}$ have solutions only for relatively big values of $r$. For $T_{\times}$ approaching the upper bound the length of the corresponding curve shrinks to a point. In many models the allowed region in the ($n_{s}$, $r$) plane has only some partial overlap with the region corresponding to the often used condition $50\leq N_{k}\leq 60$. Clearly, the approach based on well motivated bounds on $T_{\times}$ may give significantly different results compared to an (essentially arbitrary) ansatz for $N_{k}$. Thus, in the following the dependence of the results on the reheating temperature (and not on $N_{k}$) will be investigated.
@@
Results for $r\geq 10^{-4}$ are presented in figures 3 and 4 where curves of constant $T_{\times}$ are shown together with present experimental bounds and sensitivity reaches of future experiments. The present $2\sigma$ contours of the Planck+BK18 data differ quite strongly from the global fit including also ACT and DESI data (for recent discussion see e.g. [44]). The latter fit has more discriminative power for the considered models. It shows some preference for P-model with $n=1$ and for hight reheat temperature $T_{\times}$. It can oslo be reconciled with E and T-models with large values of $n$ and low reheat temperature. Actually, we see that in the considered models there is the upper bound on $n_{s}$ around 0.97, to be compared with the $2\sigma$ lower bound of the global fit at $\sim 0.965$. One can also ask how constraining are the present CMB data for the reheating temperature. Given the large difference between the Planck data and the global fit one concludes that the whole range of temperatures is acceptable. Taken Planck and the global fit separately, some constraints on the reheating temperatures can be seen in the figures. Furthermore, it is worth noting the change of ordering of the $T_{\times}={\rm const}$ curves from small to large values, as a function of $n_{s}$. This is related to the fact that $w<\frac{1}{3}$ for $n=1$ while $w>\frac{1}{3}$ for $n>2$. The case of $n=2$ is a special one, where $w=\frac{1}{3}$ and the function $r(n_{s})$ does not depend on $T_{\times}$ (the black curve in figure 3). Improving the precision of $n_{s}$ measurements, as expected for the CMB S4 experiment, would be a stronger test of the models. The expected $2\sigma$ range (0.964, 0.974) would be particularly constraining if, in addition, the upper bound on $r$ was pushed down, say to $r<10^{-5}$ (see figure 2). It also follows from figure 2 that, if the present $2\sigma$ lower bound is valid for any value of $r$, then E- and T-models with $n=1$ predict a lower bound on $r$ of the order of $10^{-10}$. The corresponding bounds for P-models and all considered models with big values of $n$ are much weaker (e.g. of the order of $10^{-26}$ for $n=10$). Some differences in the predictions for $r(n_{s})$ for a fixed reheating temperature $T_{\times}$, particularly between P-model compared to E- and T-models, can be traced back to the shape of the inflaton potential in those models.
@@ 3 $T_{\times}$ sensitive dark matter production
Many freeze-in models with the DM production described by dim 5 and dim 6 effective operators have been discussed in the literature. They include, for instance, two scalar thermal bath particles $\Phi$ annihilating into a pair of DM fermions $\chi$, $\mathcal{L}=|\Phi|^{2}\bar{\chi}\chi/\Lambda_{\rm DM}$, or two thermal bath fermions $\Psi$ annihilating either into a pair of scalar DM particles $S$, ${\cal L}={\bar{\Psi}}\Psi SS/\Lambda_{\rm DM}$ or into a pair of DM fermions $\chi$, ${\cal L}={\bar{\Psi}}\Psi{\bar{\chi}}\chi/\Lambda_{\rm DM}^{2}$.
@@ 4 Conclusions
The present and future experimental programmes will explore the properties of the Cosmic Microwave Background in a great detail. In this paper we have proposed a systematic approach to use the CMB observables and the model-independent bounds on the reheating temperature a) for discriminating between different models of inflation, b) to constrain a reheating temperature dependent particle physics scenario, freeze-in DM production models. In the first step of this approach, all independent parameters of an inflationary model are expressed in terms of the CMB observables. Next, adopting a one parameter effective description (by an effective dissipation rate $\Gamma$), a consistency relation fixes the reheating temperature in terms of the measured values of the CMB observables. For a given model of inflation the bounds on those observables then follow from model-independent bounds on the reheating temperature (see section 2). The striking consequence of those bounds is that a given model of inflation gives very narrow ranges of the power spectrum index $n_{s}$, different for different inflationary models and with very interesting dependence on the value of $r$. In this paper, the above described strategy has been applied to $\alpha$-attractor inflaton potentials, with three free parameters. We have found that the ACT data have more discriminative power than the Planck data for the considered models. They show preference for the polynomial $\alpha$-attractor P-model with $n=1$ and high reheat temperature, corresponding to $N_{k}\approx 55$. [^15] We thank Renata Kallosh and Andrei Linde for the dicussion on this point. We stress that those results are obtained for a very simple reheating model, which strongly correlates the effective equation of state parameter $w$ during the reheating period with the inflaton potential (see eq. (2.22)) .
@@ Acknowledgments
The authors would like to thank Ranata Kallosh and Andrei Linde for many discussions. A.G. thanks S K Jeesun for comments. P.K. thanks Adeela Afzal and Shiladitya Porey for sharing their Mathematica codes used to produce the experimental contours in section 2.
@@ Appendix B Simplifying assumptions underlying this analysis
  • In the calculations regarding DM we have assumed instant and local thermal equilibrium to have been reached during the reheating period. We have not considered detailed thermalisation processes like those in refs. [72, 73].